In a tennis match, with the score at deuce, the game is won by the first player who gets a clear lead of two points.If the probability that given player wins a particular point isθ, and all points are played independently, what is the probability that player eventually wins the game


1
Expert's answer
2021-04-15T07:25:54-0400

Solution:

Let us define

- sample space "\Omega" to consist of all possible infinite sequences of points;

- event "W_{i}" - nominated player wins the i th point;

- event "V_{i}" - nominated player wins the game on the i th point;

- event V - nominated player wins the game.

We want P(V). The events "\left\{W_{1}, W_{1}^{\prime}\right\}" partition "\Omega" , and thus, by the Theorem of Total Probability

"P(V)=P\left(V \mid W_{1}\right) P\left(W_{1}\right)+P\left(V \mid W_{1}^{\prime}\right) P\left(W_{1}^{\prime}\right)" ...(1)

Now "P\left(W_{1}\right)=\theta" and "P\left(W_{1}^{\prime}\right)=1-\theta" . To get "P\left(V \mid W_{1}\right)" and "P\left(V \mid W_{1}^{\prime}\right)" , we need to further condition on the result of the second point, and again use the Theorem: for example

"\begin{aligned}

P\left(V \mid W_{1}\right) &=P\left(V \mid W_{1} \cap W_{2}\right) P\left(W_{2} \mid W_{1}\right)+P\left(V \mid W_{1} \cap W_{2}^{\prime}\right) P\left(W_{2}^{\prime} \mid W_{1}\right) \\

P\left(V \mid W_{1}^{\prime}\right) &=P\left(V \mid W_{1}^{\prime} \cap W_{2}\right) P\left(W_{2} \mid W_{1}^{\prime}\right)+P\left(V \mid W_{1}^{\prime} \cap W_{2}^{\prime}\right) P\left(W_{2}^{\prime} \mid W_{1}^{\prime}\right)

\end{aligned}"

...(2)

where

"\begin{array}{ll}
P\left(V \mid W_{1} \cap W_{2}\right)=1, & P\left(W_{2} \mid W_{1}\right)=P\left(W_{2}\right)=\theta \\
P\left(V \mid W_{1} \cap W_{2}^{\prime}\right)=P(V), & P\left(W_{2}^{\prime} \mid W_{1}\right)=P\left(W_{2}^{\prime}\right)=1-\theta, \\
P\left(V \mid W_{1}^{\prime} \cap W_{2}\right)=P(V), & P\left(W_{2} \mid W_{1}^{\prime}\right)=P\left(W_{2}\right)=\theta \\
P\left(V \mid W_{1}^{\prime} \cap W_{2}^{\prime}\right)=0, & P\left(W_{2}^{\prime} \mid W_{1}^{\prime}\right)=P\left(W_{2}^{\prime}\right)=1-\theta,
\end{array}"

as,

given "W_{1} \cap W_{2}" : the game is over, and the nominated player has won given "W_{1} \cap W_{2}^{\prime}" : the game is back at deuce given "W_{1}^{\prime} \cap W_{2}" : the game is back at deuce given "W_{1}^{\prime} \cap W_{2}^{\prime}" } : the game is over, and the player has lost

and the results of successive points are independent. Thus

"\begin{aligned}

P\left(V \mid W_{1}\right) &=(1 \times \theta)+(P(V) \times(1-\theta))=\theta+(1-\theta) P(V) \\

P\left(V \mid W_{1}^{\prime}\right) &=(P(V) \times \theta)+0 \times(1-\theta)=\theta P(V)

\end{aligned}"

Hence, combining (1) and (2), we get

"\begin{array}{c}

P(V)=(\theta+(1-\theta) P(V)) \theta+\theta P(V)(1-\theta)=\theta^{2}+2 \theta(1-\theta) P(V) \\
\end{array}"

"\Rightarrow P(V)=\frac{\theta^{2}}{1-2 \theta(1-\theta)}"